OpenStudy (anonymous):

help with compound interest! please How long would it take $3,300 to grow to $9,100 if the annual rate is 4.4% and interest in compounded monthly

OpenStudy (anonymous):

\[9,100=3,300(1+\frac{.044}{12})^{12t}\] solve for \(t\)

OpenStudy (anonymous):

a) divide by 3300 first b) use the change of base formula c) divide by 12

OpenStudy (anonymous):

do you divide both sides by 3300 or no

OpenStudy (anonymous):

yes both sides

OpenStudy (anonymous):

\[\frac{91}{33}=(1+\frac{.044}{12})^{12t}\]

OpenStudy (anonymous):

then use the fact that \(A+b^x\iff x=\frac{\ln(A)}{\ln(b)}\) to find \[12t=\frac{\ln(\frac{91}{33})}{\ln(1+\frac{.044}{12})}\]

OpenStudy (anonymous):

sorry, i meant \[A=b^x\iff x=\frac{\ln(A)}{\ln(b)}\]

OpenStudy (anonymous):

last step is to divide by 12 and you are done these questions really all over the map. on line i am guessing.

OpenStudy (anonymous):

23.1 that is rounded... yes online :/

OpenStudy (anonymous):

i didn't do it, let me check

OpenStudy (anonymous):

yeah that is what i get

OpenStudy (anonymous):

cool, if you have time could we try one more? it is slightly different .... If $6,000 is placed in an account with an annual interest rate of 6%, how long will it take the amount to quadruple if the interest is compounded annually? Round your answer to two decimal places